OpenStudy (anonymous):

Solve for x: 4 over 5 x + 4 over 3 = 2x

OpenStudy (shinalcantara):

\[\frac{ 4 }{ 5x } + \frac{ 4 }{ 3 } = 2x\] is that the equation?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

@shinalcantara

OpenStudy (shinalcantara):

so finding the LCD of 5x and 3, you'll have \[\frac{ 3(4) +4(5x) }{ 15x } = 2x\] now you cross multiply

OpenStudy (anonymous):

and how do i do that??

OpenStudy (anonymous):

@shin

OpenStudy (anonymous):

@shinalcantara

OpenStudy (shinalcantara):

you multiply 15x to the right leaving you with 3(4) + 4(5x) at the left

OpenStudy (shinalcantara):

\[3(4) + 4(5x) = 15x(2x)\]

OpenStudy (shinalcantara):

you simplify and transpose everything to the right so that \[x^{2}\] will have a positive coefficient.. if you remember quadratic equation, you proceed with that

OpenStudy (shinalcantara):

*quadratic equation formula

OpenStudy (anonymous):

???????????????????

OpenStudy (anonymous):

@shinalcantara

OpenStudy (shinalcantara):

\[12x +20x = 30x ^{2}\] rearranging it you'll have: \[30x ^{2} - 20x -12 =0\] using quadratic formula: \[x = \frac{ -b \pm \sqrt{b^{2}-4ac} }{ 2a }\]

OpenStudy (shinalcantara):

a=30, b=-20, c=-12 now you substitute

OpenStudy (anonymous):

so is -b = 20? @shinalcantara

OpenStudy (shinalcantara):

yep